OpenStudy (anonymous):

Solve system of equations by elimination Solve x + 3y = 9 3x - 3y = -13

OpenStudy (anonymous):

Since the coefficient on the y-variables are opposites, we make linear combination of the two equations x + 3y = 9 3x - 3y = -13 ========= 4x =-4 x=-1 Now sub x=-1 into either of the given equations to find y -1+3y=9 3y=10 y=10/3 Solution may be written (10/3, -1) or simple as x=10/3, y=10/3